40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1541     Entry time: Sun May 3 22:48:12 2009
Author: Yoichi 
Type: Update 
Category: Locking 
Subject: Some measurements at the lock point 
I attached some measurement results at when the IFO is at the full lock point.

The first plot shows the trend of the arm powers after the interferometer was locked.
The arm powers slowly increased after the lock. This increase is observed every time the IFO is locked.
Probably this is some sort of a thermal effect (mirror lensing, PD efficiency etc).

The second plot is a CARM offset sweep. Even after the demodulation phase optimization, the lock point is not exactly at the resonance.

The third plot is the open loop TF of the AO path. The CM loop UGF is about 20kHz.
The boost and the superboost1 were turned on when this TF was measured. The IFO loses lock if the superboost2 is turned on.

TO DO LIST
Measured the DARM loop shape.
I could not turn on the dewhitening filter for ETMY. ETMX had no trouble. I will check the dewhitening circuit.
Attachment 1: ArmPowerTrend.png  12 kB  | Show | Hide all | Show all
Attachment 2: CARMSweep.png  8 kB  | Hide | Hide all | Show all
CARMSweep.png
Attachment 3: CM-AO-Loop-SB1.png  57 kB  | Hide | Hide all | Show all
CM-AO-Loop-SB1.png
ELOG V3.1.3-